# A Resolution Independent Neural Operator

### v1

@@ 4 Conclusion
In this work, we propose a resolution-independent neural operator (RINO) method that modifies the DeepONet formulation with minimal architectural changes to directly work on (rich enough) point clouds sampled from input and output functions without any underlying structural assumptions on the data structure. Specifically, we introduce a dictionary learning algorithm to identify appropriate continuous and fully differentiable basis functions for input function realizations. This allows us to construct an appropriate coordinate system to project input function point clouds onto, and use the resulting coordinates as the input function representation, which can be subsequently used in the vanilla DeepONet. The application of the introduced algorithm extends beyond the operator learning task and can be used for reconstruction tasks from incomplete or missing information, similar to GPOD. This algorithm can be considered a reconstruction method over function spaces rather than vector spaces, as is the case with GPOD or PCA.
@@ 5 Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0024162.
@@
Complementary results for numerical examples are provided in Appendices A, B and C. Appendix D lists the hyperparameters used for each problem. An ablation study for Algorithm 1, comparing our method with GPOD, is presented in Appendix E.Appendix F summarizes the formulation for Gaussian random field generations.
@@ Appendix A Nonlinear 1D Darcy’s Equation
Figure 9: (a) Loss function during optimization iterations for dictionary learning of the input function bases in the nonlinear Darcy (1D) example. (b) Distribution of reconstruction errors for the train and test datasets after training.
Figure 9(a) shows the reconstruction errors during the dictionary learning iterations. With a total of 10 learned basis functions, the data can be reconstructed fairly accurately, as shown by the distribution of errors in Figure 9(b) for both training and testing datasets.
Figure 10: (a) Loss function during optimization iterations for learning the operator in the nonlinear Darcy (1D) example. (b) Distribution of output function prediction errors for the training and testing datasets after training.
The history of the loss function during the operator learning stage is plotted in Figure 10(a). In Figure 10(b), the distribution of output function prediction errors after training is shown. The results suggest that the training is successful and the output function prediction via RINO is relatively accurate.
@@ Appendix B Nonlinear 2D Darcy’s Equation
Figure 11: (a) Loss function during optimization iterations for dictionary learning of the input function bases in the nonlinear Darcy (2D) example. (b) Distribution of reconstruction errors for the train and test datasets after training.
Figure 11(a) shows the reconstruction errors during the dictionary learning iterations. With a total of 57 learned basis functions, the data can be reconstructed fairly accurately, as shown by the distribution of errors in Figure 11(b) for both training and testing datasets. Notice that the found dictionary allows us to represent data with 57 dimensions, which is much less than the ambient dimension of the discretized input function on a $20\times 20$ grid. Therefore, we not only bypass the resolution dependence but also achieve a rich representation with lower dimensions than the ambient data dimensionality. From a computational perspective, the input dimension for the branch network is 57, whereas using an MLP branch for the vanilla DeepONet requires 400 dimensions.
Figure 12: (a) Loss function during optimization iterations for learning the operator in the nonlinear Darcy (2D) example. (b) Distribution of output function prediction errors for the training and testing datasets after training.
@@
Two examples of the test predictions for the output field are shown in Figures 13 and 14, corresponding to the 75th percentile error and the median error, respectively. In each example, the queried point cloud input function data has different numbers of points and locations because the masked information varies among them.
Figure 13: The case with output function prediction error worse than 75% of the test data. For the arrangement description of the subfigures, refer to Figure 7.
Figure 14: The case with output function prediction error worse than half of the test data. For the arrangement description of the subfigures, refer to Figure 7.
@@ Appendix C Burgers’ Equation
Figure 15(a) shows the reconstruction errors during the dictionary learning iterations. The distribution of reconstruction errors in Figure 15(b) for both training and testing datasets indicates that the found dictionary can satisfactorily represent the input function data.
Figure 15: (a) Loss function during optimization iterations for dictionary learning of the input function bases in Burger’s example. (b) Distribution of reconstruction errors for the train and test datasets after training.
Figure 16(a) illustrates the evolution of the loss function during the operator learning stage. The distribution of output function prediction errors after training is depicted in Figure 16(b). These results indicate that the training was successful, leading to relatively accurate output function predictions by RINO.
Figure 16: (a) Loss function during optimization iterations for learning the operator in the Burger’s example. (b) Distribution of output function prediction errors for the training and testing datasets after training.
Appendix D Network Architecture
@@
The hyperparaemtrs utilized for learning the dictionary are listed in Table LABEL:tab:hyper-dl.
@@
Antiderivative Nonlinear 1D Darcy Nonlinear 2D Darcy Burger’s
Number of hidden units $20$ $30$ $50$ $30$
Number of hidden layers $2$ $2$ $3$ $2$
Learning rate $1.66\times 10^{-4}$ $3\times 10^{-4}$ $6.6\times 10^{-4}$ $6.6\times 10^{-4}$
$\lambda$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-5}$
$\omega_{0}$ $5$ $10$ $10$ $10$
@@
Appendix E Ablation Study: Comparison with Gappy Proper Orthogonal Decomposition
@@
Note that the basis functions used in the generative function, Equation (21), are purposely chosen to not be orthogonal (see Figure 19(a)) and exhibit a multiscale nature, as they have different length scales. It may be of interest to investigate whether our algorithm can find orthogonal basis functions, how similar these basis functions are to those found by GPOD, and how closely they resemble the non-orthogonal basis functions used for generating the data. It is known that the basis functions learned by PCA-based algorithms, such as GPOD, will be orthogonal.
@@
Figure 19: Orthogonality check among basis functions: (a) Basis functions used to generate random function data (see Equation 21). (b) Basis vectors identified by the Gappy POD algorithm, and (c) basis functions identified by the proposed algorithm with 50% of the data masked.
@@
Appendix F Random Field Generation

### v2

@@ 4 Conclusion
In this work, we propose a resolution-independent neural operator (RINO) that modifies the DeepONet formulation with minimal architectural changes to directly work on (sufficiently rich) point clouds sampled from input and output functions without any underlying structural assumptions on the data structure. Specifically, we introduce a dictionary learning algorithm to identify appropriate continuous and fully differentiable basis functions for input function realizations. This allows us to construct an appropriate coordinate system (frame) to project input function point clouds onto, and use the resulting coordinates as the input function representation, which can be subsequently used in the vanilla DeepONet. The application of the introduced algorithm extends beyond the operator learning task and can be used for reconstruction tasks from incomplete or missing information, similar to GPOD. This algorithm can be considered a reconstruction method over function spaces rather than vector spaces, as is the case with GPOD or PCA.
@@ 5 Acknowledgements
We acknowledge the discussion with Professor George E. Karniadakis during the CRUNCH seminar, who brought random projection and ReNO [3] to our attention, which improved our work. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0024162.
@@
Complementary results for the numerical examples are provided in Appendices A, B and C. Appendix E lists the hyperparameters used for each problem. An ablation study for Algorithm 1, comparing our method with GPOD, is presented in Appendix F. Moreover, the use of random projections as an alternative for finding vector representations of input or output function data has been studied and compared to our proposed dictionary learning algorithm. Appendix G summarizes the formulation for Gaussian random field generation.
@@ Appendix A Nonlinear 1D Darcy’s Equation
Figure 14(a) shows the reconstruction errors during the dictionary learning iterations. With a total of 10 learned basis functions, the data can be reconstructed with high accuracy, as shown by the distribution of errors in Figure 14(b) for both training and testing datasets.
Figure 14: 1D Nonlinear Darcy Example: (a) Loss function during optimization iterations for dictionary learning of the input basis functions. (b) Distribution of reconstruction errors for the train and test datasets after training.
The history of the loss function during the operator learning stage is plotted in Figure 15(a). In Figure 15(b), the distribution of output function prediction errors after training is shown. The results suggest that the training is successful and the output function prediction via RINO is very accurate.
Figure 15: 1D Nonlinear Darcy Example: (a) Loss function during optimization iterations for operator learning. (b) Distribution of output function prediction errors for the training and testing datasets after training.
@@ Appendix B Nonlinear 2D Darcy’s Equation
Figure 16(a) shows the reconstruction errors during the dictionary learning iterations. With a total of 57 learned basis functions, the data can be reconstructed with high accuracy, as shown by the distribution of errors in Figure 16(b) for both training and testing datasets. Notice that the learned dictionary allows us to represent data with 57 dimensions, which is much less than the ambient dimension of the discretized input function on a $20\times 20$ grid. Therefore, we not only bypass the resolution dependence but also achieve a rich representation with lower dimension than the ambient data dimensionality. From a computational perspective, the input dimension for the branch network is 57, whereas using an MLP branch for the vanilla DeepONet requires 400 dimensions.
Figure 16: 2D Nonlinear Darcy Example: (a) Loss function during optimization iterations for dictionary learning of the input basis functions. (b) Distribution of reconstruction errors for the train and test datasets after training.
@@
Figure 17: 2D Nonlinear Darcy Example: (a) Loss function during optimization iterations for operator learning. (b) Distribution of output function prediction errors for the training and testing datasets after training.
Two examples of the test predictions for the output field are shown in Figures 18 and 19, corresponding to the 75th percentile error and the median error, respectively. In each example, the queried point cloud input function data has different numbers of points and locations because the masked information varies among them.
Figure 18: 2D Nonlinear Darcy Example: The case with 75th percentile output function prediction error. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
Figure 19: 2D Nonlinear Darcy Example: The case with 50th percentile output function prediction error. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
@@ Appendix C Burgers’ Equation
Figure 20(a) shows the reconstruction errors during the dictionary learning iterations. The distribution of reconstruction errors in Figure 20(b) for both training and testing datasets indicates that the learned dictionary, with a total of 16 basis functions, can satisfactorily represent the input function data.
Figure 20: Burgers’ Equation: (a) Loss function during optimization iterations for dictionary learning of the input basis functions. (b) Distribution of reconstruction errors for the train and test datasets after training.
Figure 21(a) illustrates the evolution of the loss function during the operator learning stage. The distribution of output function prediction errors after training is depicted in Figure 21(b). These results indicate that the training was successful, leading to accurate output function predictions by RINO.
Figure 21: Burgers’ Equation: (a) Loss function during optimization iterations for operator learning. (b) Distribution of output function prediction errors for the training and testing datasets after training.
Appendix D Nonlinear 2D Darcy’s Equation on Unstructured Meshes
Figure 22(a) shows the reconstruction errors during the dictionary learning iterations for the input function point cloud data. The distribution of reconstruction errors in Figure 22(b) suggests that a satisfactory reconstruction error can be achieved by the learned dictionary with 40 basis functions.
Figure 22: 2D Darcy’s Equation on Unstructured Meshes: (a) Loss function during optimization iterations for dictionary learning of the input basis functions. (b) Distribution of reconstruction errors for the train and test datasets after training.
The evolution of the loss function during the operator learning stage is depicted in Figure 23(a) where both branch and trunk nets are unknown a priori. The distribution of output function prediction errors after training is plotted in Figure 23(b). A successful training with satisfactory accuracy in RINO predictions is confirmed by the results.
Figure 23: 2D Darcy’s Equation on Unstructured Meshes: (a) Loss function during optimization iterations for operator learning. (b) Distribution of output function prediction errors for the training and testing datasets after training.
As discussed in the main text, the proposed method can flexibly use a predefined trunk network by substituting a learned dictionary of approporiate basis functions for the output function data. Figure 24 shows the reconstruction error distributions for the output function point cloud data, where Algorithm 2 was able to identify an appropriate dictionary of 54 basis functions that span the output solution fields.
Figure 24: 2D Darcy’s Equation on Unstructured Meshes: Distribution of reconstruction errors for the training and test datasets after learning the dictionary of basis functions for the output function.
The evolution of the loss function during the operator learning stage for the case where we utilized the learned, predefined trunk dictionary is depicted in Figure 25(a), where only the branch network is trained. The distribution of output function prediction errors after training is plotted in Figure 25(b). A successful training with satisfactory accuracy is confirmed by the results for the case where we indirectly learned a mapping between the coefficients of two sets of weakly orthogonal basis functions: one corresponding to the input function data and the other to the output function data, which were identified separately and offline.
Figure 25: 2D Darcy’s Equation on Unstructured Meshes: (a) Loss function during optimization iterations for operator learning with the predefined dictionary learned trunk net. (b) Distribution of output function prediction errors for the training and testing datasets after training.
Appendix E Network Architecture
@@
The hyperparaemtrs utilized for learning the dictionary are listed in Table LABEL:tab:hyper-dl. The column “2D Darcy*” corresponds to the last numerical example, where we solve the nonlinear 2D Darcy problem on different finite element meshes from one realization to another.
@@
Antiderivative 1D Darcy 2D Darcy Burger’s 2D Darcy*
Number of hidden units $20$ $30$ $50$ $30$ $20$
Number of hidden layers $2$ $2$ $3$ $2$ $2$
Learning rate $1.66\times 10^{-4}$ $3\times 10^{-4}$ $6.6\times 10^{-4}$ $6.6\times 10^{-4}$ $4\times 10^{-4}$
$\lambda$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-5}$ $10^{-3}$
$\omega_{0}$ $5$ $10$ $10$ $10$ $10$
@@
  • Nonlinear 2D Darcy*: SIREN[100,100,100] with $w_{0}=10$
For the “nonlinear 2D Darcy*” case, we also compare it with the scenario where our dictionary learning algorithm is applied to output function realizations to learn output basis functions offline and use them as a predefined trunk net. In this setup, each output basis was parametrized by SIREN[20|20] with $w_{0}=10$. The sparsity parameter $\lambda=10^{-4}$ and the learning rate is $10^{-3}$.
Appendix F Ablation Study: Comparison with Gappy Proper Orthogonal Decomposition and Random Projection
@@
Note that the basis functions used in the generative function, Eq. (24), are purposely chosen to not be orthogonal (see Figure 28(a)) and exhibit a multiscale nature, as they have different length scales. It may be of interest to investigate whether our algorithm can find orthogonal basis functions, how similar these basis functions are to those found by GPOD, and how closely they resemble the non-orthogonal basis functions used for generating the data. It is known that the basis functions learned by PCA-based algorithms, such as GPOD, will be orthogonal.
@@
Figure 28: Orthogonality check among basis functions: (a) Basis functions used to generate random function data (see Eq. 24). (b) Basis vectors identified by the Gappy POD algorithm, and (c) basis functions identified by the proposed algorithm with 50% of the data masked.
@@
One may consider using random smooth functions, such as those parameterized by a shallow neural network [33, 31, 32, 16], as a method to construct basis functions in a predefined manner. These functions can naturally impose smoothness as inductive biases and can remain resolution-independent during the reconstruction and projection steps. Here we aim to benchmark and compare their performance using this pedagogical example. To this end, we consider two cases: 1) random shallow ReLU networks, where each basis is a ReLU network with 20 hidden neurons randomly initialized, and 2) classical cosine basis functions with random frequency and phase shifts as follows:
$\psi_{i}(x)=\cos(w_{i}x+b_{i});w_{i}\sim\mathcal{N}(0,\sigma=2\pi),b_{i}\sim\text{Uniform}[0,2\pi],$ (25)
for the $i$-th cosine basis functions. For completeness, we also consider the classical monomial and Legendre basis functions (deterministic). In Figure 30, we compare the reconstruction performance of these sets of basis functions by increasing the number of basis functions used in each set and by gradually sparsifying the data.
When the data density is sufficiently high, as shown in Figure 30(a), random ReLU, random cosine, and Legendre bases deliver satisfactory performance. However, in our numerical experiments, they appear to require two orders of magnitude more basis functions than ours, which is not ideal for downstream learning tasks, as one would prefer to operate with an embedding space of dimension 3 rather than 100. This preference is due to both computational efficiency and the fact that the amount of data required for such learning tasks may scale exponentially with respect to dimensionality. As data sparsity increases, as shown in Figure 30(b and c), the performance gap between our method and others grows more pronounced. For instance, with 85% of the data masked in Figure 30(c), the performance of the other methods becomes unacceptable. This decline is due to the limited data, making these other methods more prone to overfitting.
In all cases, monomial basis functions did not perform well, regardless of the number of chosen basis elements or the amount of data available. This may be due to the fact that monomial bases of type $\psi_{n}(x)=x^{n}$ lack the appropriate inductive biases needed to effectively learn the data, particularly when the data exhibits sinusoidal behavior. This observation underscores the importance of selecting a sufficiently flexible set of basis functions, which is often not obvious at first. Therefore, an approach that can automatically identify the right set of bases in a data-driven manner is crucial, and this is one of the key contributions of this paper.
(a) 25% Masked
(b) 50% Masked
(c) 85% Masked
Figure 30: Comparison of reconstruction accuracy on test data using four distinct sets of basis functions. Each set is evaluated through five independent computations, with randomly chosen parameters for stochastic methods; for deterministic basis functions (monomial and Legendre), no randomness is involved. As the results demonstrate, even as data sparsity increases, our method consistently outperforms alternatives. Notably, our approach achieves this with significantly fewer basis functions, utilizing only 3 compared to 100 in other methods.
To further analyze the performance of random bases in the case where the data has the highest sparsity (i.e., 85% masked data), we plotted the reconstruction curves for the worst case and the 25th quintile in terms of relative errors in Figure 31. As expected (and as consistent with the literature [31, 16]), the random basis functions perform well in the interpolation regime but poorly in the extrapolation regime beyond the data support (as observed in the purple-shaded areas in these figures). Moreover, the underlying generative process of the data is symmetric at $x=0$, see Figure 26(a). However, there is no guarantee that this symmetry can be recovered using random basis functions. As already discussed, these issues do not arise with our method, as we learn appropriate basis functions by considering the common correlation structure across all realizations, rather than merely interpolating for each individual realization.
(a) Random Projection via ReLU Basis function
(b) Random Projection via Cosine Basis Functions
Figure 31: Reconstruction curves for the worst case and the 25th quantile in terms of relative reconstruction error in the case where 85% of the data was masked.
Spanning versus Parametrization
As we have shown, linear (with respect to the basis functions) projection and reconstruction using linearly independent random basis functions often requires exponentially more basis functions compared to an optimal orthogonal basis set, provided the signals of interest can be spanned by such an optimal orthogonal basis. Although this is aligned with the linear theory [16], it does not take advantage of an additional possibility, consistent with established embedding theorems, which suggests that one can embed an $n$-dimensional manifold with at most $12n+$ 1 generic observables [43, 44] (in our case, the components of our data in $12n+$ 1 random basis functions). These projections do not span our intrinsically $n$-dimesional data; yet, the projections of the data in any additional random basis functions can be obtained as functions of the first $12n+$ 1 ones. We say that the $12n+$ 1 basis functions parametrize the data, but do not span them. The approximation of the data can now be achieved via nonlinear parameterization, which maps the projected solution on the hyperplane back to the target manifold. This approach shares similarities with the post-processing Galerkin method [20, 14]. However, such nonlinear parameterization may have some caveats: (1) it is less interpretable than our proposed spanning scheme; (2) the number of basis functions is not known a priori—only an upper bound is provided—while our approach adaptively identifies the number of basis functions according to a tolerance; (3) The embedding space constructed via random projection is highly sensitive to sampling, which undermines the uniqueness of the signal representation. To elaborate further on the last point, we conducted a simple test as follows. We pick one of the signal realization from the test data set and aim to reconstruct the signal with three linearly independent random cosine basis, as mentioned earlier. To assess the sensitivity of the embeddings, we consider five cases where half of the points are masked in each case. However, the locations of these masked points are randomly selected and vary from one case to another. The embeddings obtained from random projection, shown in the right column of Figure 32(a), are clearly dependent on the location of the samples. The left column shows the reconstructed signal, where we again observe high variance. In contrast, the projection via optimal orthogonal basis functions, as shown in Figure 32(b),is consistent across cases and independent of the sample locations.
(a) Random Basis
(b) Optimal Basis
Figure 32: Consistency check of the embeddings obtained from projections via (a) random basis and (b) optimal basis sets. The right column display the coefficients corresponding to each basis function, while the left column show the reconstructed full fields based on the identified coefficients. Different colors represent distinct queried datasets, where the number of data points remains the same but their locations are chosen randomly and differ from one dataset to another.
Appendix G Random Field Generation